OpenStudy (anonymous):

Solve for y. M=2x+3y

OpenStudy (zarkon):

subtract 2x divide by 3

OpenStudy (rosey):

zarkon help meee

OpenStudy (zarkon):

don't you have people helping you...looks like it :)

OpenStudy (anonymous):

so it would be... m-2x over 3 = y?

OpenStudy (zarkon):

\[\frac{M-2x}{3}=y\] ues

OpenStudy (zarkon):

yes